OpenStudy (anonymous):

The acceleration due to gravity on the moon is 1.62 m/s^2, approximately a sixth of the value on earth. For a given velocity V and a given angle theta, the ratio of the range of an ideal projectile on the moon to the range of the same projectile on earth Rmoon/Rearth will be?

OpenStudy (anonymous):

since nobody helped with the last one lol...figured try this one

OpenStudy (anonymous):

The answer is 1/6.. Is this physics or math?

OpenStudy (anonymous):

physics...i put in both sections

OpenStudy (anonymous):

it's multiple choice.... a. 2 b. 4 c. 6 d. 3 and e is 12

OpenStudy (anonymous):

so...don't think itll be 1/6

OpenStudy (anonymous):

This was how I did it... so projectile motion is just horizontal motion, and the formula for x (the distance) = 1/2at^2... If I divide Rmoon by Rearth = the time cancels out so it will be \[\left(\begin{matrix}1/2(1.62) \\ 1/2(9.72)\end{matrix}\right)\] = 1/6......

OpenStudy (anonymous):

it's multiple choice.... a. 2 b. 4 c. 6 d. 3 and e is 12

OpenStudy (anonymous):

Yeah, so I don't know... I was just showing you in case you wanted to know how I got my answer...

OpenStudy (anonymous):

if i actually just solve that i get 4....so i dont know now